OpenStudy (anonymous):

Doug is 23 years older than Danielle. In 11 years, he will be twice as old as Danielle. How old is each of them now?

OpenStudy (callisto):

let danielle be x years old now, doug is x+23 years old 2(x+11 ) = x+23+11 x=12 therefore danielle is 12 years old now and doug is 35 years old now

OpenStudy (anonymous):

Thank you!